what is The pH and pOH of a 1.00x10-3M solution of CH3COOH (Ka=1.75x10-5) ?
1
Expert's answer
2020-09-22T03:37:03-0400

Solution:

"CH_{3}COOH=CH_{3}COO^{-}+H^{+}"


"K_{a}=1.75*10^{-5}" so this equilibrium lies mostly to the left and there is no need for ICE tables. Under these conditions the equilibrium expression simplifies to 


"[H^{+}]= \sqrt{\smash[b]{K_{a}*C_{M}}}"


"[H^{+}]= \sqrt{\smash[b]{(1.75*10^{-5})*(10^{-3})}}"


"[H^{+}]=" "1.32*10^{-4} M"



"pH=-log(1.32*10^{-4})"


"pH=3.88"


"pOH=14-pH=10.12"